OpenStudy (baldymcgee6):

Without using L’Hôpital’s Rule, evaluate the following limit. Use ∞,−∞ to denote divergence to +∞,−∞ and DNE to indicate an other type of non-convergence.

OpenStudy (baldymcgee6):

\[\lim_{x \rightarrow \infty} (2\sin(lnx) +1)/x^2\]

OpenStudy (anonymous):

Funny really .... the denumerator is bounded.... and the denumerator goes to infinity. Any questions ?!

OpenStudy (anonymous):

numerator is bounded, denominator is unbounded.... so limit is zero?

OpenStudy (anonymous):

Are you the 99% that own the 1% OR the 1% that owns the 99 ?

OpenStudy (baldymcgee6):

@Mikael I have no idea what you are talking about

OpenStudy (anonymous):

basics of limits\[\frac{ bounded-function }{ \infty } = 0\]

OpenStudy (baldymcgee6):

So how do I do my problem? haha

OpenStudy (anonymous):

sine is stuck between -1 and 1, so numerator is stuck between -1 and 3

OpenStudy (anonymous):

I don't even see how L'hopital is applicable here?

OpenStudy (anonymous):

it isn't

OpenStudy (anonymous):

You write: As a we have learned/appears-on-page{insert num. p.} a limit of a a ratio of bounded function by another f. that tends to +infty is 0. Now as we know |sin x| =< 1 so it applies.

OpenStudy (anonymous):

the largest the numerator can be is 3 the denominator goes to infinity as fast as, well as fast as \(x^2\)

OpenStudy (baldymcgee6):

@satellite73, but we can't have a negative ln either

OpenStudy (anonymous):

you are taking the limit as \(x\to \infty\) so presumably it is positive

OpenStudy (anonymous):

By the way is it a custom to GIVE or NO-TO-GIVE medals for a complete and correct solution ?

OpenStudy (baldymcgee6):

Hold your horses mikael, ;)

OpenStudy (anonymous):

No problem amigo !

OpenStudy (baldymcgee6):

@Mikael, so how do I present my work?

OpenStudy (anonymous):

You write: "As a we have learned/appears-on-page{insert num. p.} a limit of a a ratio of bounded function by another f. that tends to +infty is 0. Now as we know |sin x| =< 1 so it applies."

OpenStudy (anonymous):

As a teacher I always would give full mark for solution that is the simplest and shortest of all the CORRECT solutions. This answer satisfies these requirements

OpenStudy (baldymcgee6):

You're a teacher?

OpenStudy (anonymous):

and more

OpenStudy (baldymcgee6):

aren't we all. . .

OpenStudy (anonymous):

well - what is it that you say ? Do you think there is some common sense in your sequence of remarks and requests ? BTW don't bother - this is a retorical question

OpenStudy (baldymcgee6):

I was simply implying that we are all teachers in our own sense. @Mikael